OpenStudy (anonymous):

so for (x-3)^2+(y+1)^2=4 h=3,k=-1 r=2 right @satellite73

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

yeah that looks right to me! Center is (h,k) Radius (r) h = 3 k = -1 r = 2 So yeah looks correct to me

OpenStudy (anonymous):

ty!!!!! and what do you think about x^2+(y-2)^2=1? r= h= k=

OpenStudy (anonymous):

would h be 0?@johnweldon1993

OpenStudy (johnweldon1993):

Yeah h would be 0....k would be 2 (so the center would be at point (0,2) on a graph... what about r...?

OpenStudy (anonymous):

r=1

OpenStudy (johnweldon1993):

Perfect :)

OpenStudy (anonymous):

U r the best thank you!

OpenStudy (johnweldon1993):

 So basically something like that lol

OpenStudy (johnweldon1993):

And no problem! :)

OpenStudy (anonymous):

yea looks right lol :)

OpenStudy (anonymous):

just one more question, what if the problem looks like 2x^2+2y^2+20x+16y+74=0?

OpenStudy (johnweldon1993):

I hate those lol...hang on let me figure that out...

OpenStudy (anonymous):

lol k

OpenStudy (johnweldon1993):

Lol it's a factoring process but I can't QUITE remember how to do it...still working on it though lol

OpenStudy (anonymous):

ok thanks no worries i am trying to as well :)

OpenStudy (anonymous):

is it 2(x+5)^2+2(y+4)^2-8=0 @johnweldon1993

OpenStudy (johnweldon1993):

-8 of course....forgot about that part XD yes that makes sense Good job! Then dividing everything by 2 makes it (x + 5)² + (y + 4)² = 4 So h = ? k = ? r = ?

OpenStudy (anonymous):

h=5 k=4 r=2

OpenStudy (johnweldon1993):

check the h and k again...

OpenStudy (anonymous):

h=0 k=2 r=1

OpenStudy (anonymous):

but I dont think I factored it correctly X(

OpenStudy (johnweldon1993):

No....with the equation 2(x+5)^2+2(y+4)^2-8=0 *you factored it correctly, I have the same thing....but you should see that h = -5 k = -4 and r = 2 Let me double check with wolfram....

OpenStudy (anonymous):

lol I check it with wolfram and got something else

OpenStudy (johnweldon1993):

Yeah wolfram has the same thing h at -5 k at -4 and the radius looks to be 2

OpenStudy (anonymous):

ok guess I gotta go with it

OpenStudy (johnweldon1993):

lmao really ? http://www.wolframalpha.com/input/?i=2x%5E2%2B2y%5E2%2B20x%2B16y%2B74%3D0%3F

OpenStudy (anonymous):

haha yup but thank you so much !!!!!!!!!!!!!!!!!!!!!!!@johnweldon1993